# Sequential Model for Predicting Patient Adherence in Subcutaneous Immunotherapy for Allergic Rhinitis

### v2

@@ Abstract
Objective: Subcutaneous Immunotherapy (SCIT) is the long-lasting causal treatment of allergic rhinitis. How to enhance the adherence of patients to maximize the benefit of allergen immunotherapy (AIT) plays a crucial role in the management of AIT. This study aims to leverage novel machine learning models to precisely predict the risk of non-adherence of patients and related systematic symptom scores, to provide a novel approach in the management of long-term AIT.
Methods: The research develops and analyzes two models, Sequential Latent Actor-Critic (SLAC) and Long Short-Term Memory (LSTM), evaluating them based on scoring and adherence prediction capabilities.
Results: Excluding the biased samples at the first time step, the predictive adherence accuracy of the SLAC models is from $60\,\%$ to $72\%$, and for LSTM models, it is $66\,\%$ to $84\,\%$, varying according to the time steps. The range of Root Mean Square Error (RMSE) for SLAC models is between $0.93$ and $2.22$, while for LSTM models it is between $1.09$ and $1.77$. Notably, these RMSEs are significantly lower than the random prediction error of $4.55$.
Conclusion: We creatively apply sequential models in the long-term management of SCIT with promising accuracy in the prediction of SCIT nonadherence in Allergic Rhinitis (AR) patients. While LSTM outperforms SLAC in adherence prediction, SLAC excels in score prediction for patients undergoing SCIT for AR. The state-action-based SLAC adds flexibility, presenting a novel and effective approach for managing long-term AIT.
@@ 2 Introduction
Allergic rhinitis is characterized by allergen-specific IgE-mediated inflammation in upper respiratory inflammation with a prevalence of up to $30\,\%$ worldwide (Meltzer, 2016). In addition to allergen avoidance as the superior criterion, allergen-specific immunotherapy (AIT) aims to induce specific allergen immune tolerance, consequently achieving a status of clinical symptom remission. The repeatable intake of the specific unmodified or chemically modified allergens (allergoids) was the key to maintaining the symptoms. Among these approaches of AIT, subcutaneous immunotherapy (SCIT), sublingual immunotherapy (SLIT), and lymphatic immunotherapy (SLIT) are demonstrated as the mainstream treatments regarding efficacy, safety, and side effects. Compared to the SLIT, SCIT is a clinic-dependent treatment in which the patient accepted an allergen extract injection subcutaneously. It is divided into the initial treatment stage (dose accumulation stage) and the maintenance treatment stage (dose maintenance stage). The World Health Organization (WHO) recommends that immunotherapy be maintained for three to five years and clinically recommended for at least two years. Patient adherence is a critical factor in ensuring long-lasting efficacy and sustaining symptom-relieving effects.
Due to the long duration of SCIT, cumbersome process, slow onset, individual differences in treatment effect, and other factors fundamentally impact the completeness of therapeutics. From the reported studies on AIT, the rate of adherence ranged from around $25\%$ to over $90\%$ (Passalacqua et al., 2013). The WHO adopted the definition of ‘adherence’ as “the extent to which a person’s behavior, such as taking medication, following a diet, or executing lifestyle changes, corresponds with agreed recommendations from a health care provider” (Organization et al., 2003). In recent EAACI guidelines, it is highlighted to educate patients on how immunotherapy works and on explaining the importance of compliance to the regular doses for three years of treatment(Roberts et al., 2018).
@@
In healthcare, machine learning, especially sequential models, stands at the forefront of innovation, providing new ways to analyze complex medical data and improve patient treatments. These models excel in processing and analyzing time-dependent data, making them ideal for predicting patient adherence to treatments like SCIT for AIT. By effectively utilizing sequential data, these algorithms uncover hidden temporal patterns and correlations, leading to more accurate and personalized treatment plans. In this study, we have selected and evaluated two specific sequential models tailored to this scenario. Our findings demonstrate that these models are not only effective in predicting patient adherence to medical treatments but also invaluable in enhancing treatment strategies, thereby making a significant contribution to patient-centered healthcare.
3 related work
Non-sequential models for adherence prediction. A hybrid model was presented to combine artificial neural networks (ANNs) and the genetic algorithm (GA) to examine 26 factors influencing adherence to dietary plans (Mousavi et al., 2022). Wang et al. (2020) developed two machine learning methods, neural network (NN) and support vector machine (SVM), to predict nonadherence in Crohn’s Disease patients. These models assist caregivers by streamlining the intervention process using azathioprine (AZA). Warren et al. (2022) investigated the relationship between patient adherence to prescribed medication for opioid use disorder and various risk factors in patients with opioid use disorder. Four different machine learning classifiers were proposed, including Logistic Regression, Decision Tree, Random Forest, and XGBoost. They compared the performance of these classifiers to explore the association. Since the adherence threshold can determine the prediction model’s performance by the binary adherence status, Ruff et al. (2019) proposed a simulation framework to assess the performance of prediction models based on different adherence thresholds.
Sequential models for adherence prediction. Hsu et al. (2022) investigated the advantages of incorporating patient history into the prediction of medication adherence. They assessed the performance of temporal deep learning models, particularly LSTM and simple Recurrent Neural Networks (RNN), and compared these with non-temporal models such as Multilayer Perceptrons (MLP), ridge classifiers, and logistic regression. To optimize the efficacy of cognitive training for older adults, Singh et al. (2022) employed multivariate time series analysis and developed personalized models for each patient to capture their unique adherence patterns. However, the sequential data of patients is often characterized by fluctuating adherence and high dropout rates, resulting in uneven, unaligned, and missing values in the time series data. To address this challenge, Schleicher et al. (2023) applied change point detection to identify phases with varying dropout rates, presented methods for handling uneven and misaligned time series, and utilized time series classification to predict the user’s phase.
@@ 4 Methods
In this section, we outline the comprehensive methodologies employed in our study. These methods are designed to rigorously analyze the effectiveness of SCIT in treating AR patients. We detail the study design, population criteria, treatment protocols, and sequential models to provide a clear understanding of our research framework.
@@ 4.1 Study design
The study design is a critical component that shapes the direction and reliability of our research. It includes a systematic approach to selecting the study population, the treatment methods applied, and the evaluation criteria.
@@ 4.1.2 SCIT treatment and evaluation
[…]
Patients receive regular treatment evaluations, including symptom scores and medication scores. The symptom score recorded a total of nasal symptoms (nasal itching, sneezing, rhinorrhea, nasal congestion), ocular symptoms (ocular itching, lacrimation), and pulmonary symptoms (shortness of breath, tightness in chest, perennial cough, wheezing), and assessed symptom severity using the visual analogue scale (VAS). In the VAS score, the score of each symptom is from 0 to 10. 0 indicates that the patient has no discomfort and 10 indicates that the patient is very uncomfortable. The patient gives the score of each symptom according to the actual situation, and the sum of all symptom scores is the symptom score. Medication score recorded the use of current adjuvant medication within 1 month to reach symptom relief. The use of oral antihistamines, antileukotrienes, and bronchodilators were recorded as one point, local glucocorticoids as two points, oral glucocorticoids or combined medication (hormones and $\beta 2$ receptor agonists) as three points, and the total cumulative score was the medication score. Symptom scores and medication scores were assessed once at registration of SCIT and then thereafter.
@@ 4.1.3 Data Collection
Data were collected from patient records in hospitals, and the following information was extracted for analysis: patient age, gender, distance to clinic, ratio of AIT cost to family income, allergen test results, etc., as well as patient VAS system score and medication score information, including baseline data of patients before injection therapy, adverse reactions to SCIT. For the descriptive analysis, categorical variables were given as numbers and percentages, and continuous variables were presented using mean, standard deviation, median, interquartile range (IQR), and minimum and maximum values.
@@ 4.1.4 Survey methods
Adherence was defined as the accomplishment of three years of AIT including the patients further received AIT. Non-adherence was defined as discontinuation of AIT at random time points during three years. The follow-up contents included (1) the main reasons for patients’ discontinuation of treatment; (2) the duration of discontinuation of treatment, and (3) Allergic symptoms after discontinuation of treatment.
@@ 4.2.1 Data
gender Female 62 (30) 22 40
Male 143 (70) 51 92
distance to clinic(km) $\leq 10$ 136 (66) 56 80
$>10$ 69 (34) 17 52
@@
Figure 1: Histogram of scores across six-time steps. Score value (horizontal axis) vs. count (vertical axis).
@@
We have a dataset $D$, comprising sequences $x_{1},\dots,x_{T}\in R^{11}$, $y_{1},\dots,y_{T-1}\in R^{1}$, and a corresponding action $a_{t}\in R^{1}$. In the context of healthcare, the observations encompass $y_{t}$ (see Table 2) whether the patient will cease the treatment in the interval between the scoring measurements at $x_{t}$ and $x_{t+1}$. The actions $a_{t}$ represent the ongoing medical procedures for the patient during the period from $x_{t}$ to $x_{t+1}$ (see Fig. 1). In this context, $a_{t}$ is binary, reflecting whether treatment is given, and is numerically equivalent to the adherence variable $y_{t}$. Despite their numerical equivalence, we maintain a distinction between action $a_{t}$ and adherence $y_{t}$ to enhance model clarity and accommodate future research expansions, potentially allowing for a wider range of action values. For each patient, we possess basic information $s\in R^{14}$ which includes age, gender, commute distance to clinic, ratio of cost to family income, eosinophils count, eosinophils percentage, nasal allergen provocation test (change of nasal resistance, $\Delta\text{NR}(\%)$), peak nasal inspiratory flow. $\Delta\mathrm{PNIF}(\%)$), serum total IgE level, sIgE of Dermatophagoides pteronyssinus (Derp), sIgE of Dermatophagoides farinae (Derf), skin prick test (Derp, Derf) (see Table 1 for more details).
4.2.2 Sequential latent variable models
[…]
figures/lvm
Figure 2: SLAC model schematic. Solid and dashed lines denote the generative and inference model pathways, respectively. The gray circles represent observed data, and the white circles denote latent variables. The figure is adapted from (Lee et al., 2020).
@@ 4.2.2 Sequential latent variable models
In our research, we employ the Stochastic Latent Actor-Critic (SLAC) model (Lee et al., 2020). However, our application differs from the original use of SLAC which is typically associated with reinforcement learning. Instead, we leverage its sequential latent-variable model without Actor-Critic. This approach aligns with similar methodologies found in other works (Krishnan et al., 2015; Karl et al., 2016; Gregor et al., 2018). The choice of the SLAC model was motivated by its ability to facilitate more efficient learning and superior generalization in intricate environments.
The sequential latent variable model consists of an inference model and a generative model (see Fig. 2). The inference model in a sequential latent variable model typically aims to approximate the posterior distribution of the latent variables given the observed data. It tries to infer the hidden states $z$ based on the observed inputs $x$ and some internal state $s$. The inference models the probability distributions of the latent variables $z^{1}$ and $z^{2}$ at different time steps. $q_{\phi}$ denotes the variational distribution parameterized by $\phi$,
$z^{1}_{1}$ $\sim q_{\phi}(z_{1}^{1}|x_{1},s)$ (1)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}|z_{1}^{1})$ (2)
$z^{1}_{t+1}$ $\sim q_{\phi}(z^{1}_{t+1}|x_{t+1},z^{2}_{t},a_{t})$ (3)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}|z^{1}_{t+1},z^{2}_{t},a_{t}).$ (4)
The generative model, on the other hand, describes how the observed data is generated from the latent variables. It defines a process by which the latent states $z$ give rise to the observations $x$. The generative model is described by a set of equations that represent the probability distributions of both the initial latent states and their transitions over time, as well as the likelihood of the observations given the latent states, with $p_{\phi}$ indicating the parameterized generative distribution.
$z^{1}_{1}$ $\sim p(z_{1}^{1})$ (5)
$z^{2}_{1}$ $\sim p_{\phi}(z^{2}_{1}|z_{1}^{1})$ (6)
$z^{1}_{t+1}$ $\sim p_{\phi}(z^{1}_{t+1}|z^{2}_{t},a_{t})$ (7)
$z^{2}_{t+1}$ $\sim p_{\phi}(z^{2}_{t+1}|z^{1}_{t+1},z^{2}_{t},a_{t})$ (8)
$x_{t}$ $\sim p_{\phi}(x_{t}|z_{t}^{1},z^{2}_{t})$ (9)
$y_{t}$ $\sim p_{\phi}(y_{t}|z_{t}^{1},z^{2}_{t}).$ (10)
We have the evidence lower bound (ELBO):
$\log p_{\phi}(x_{1:t+1}|a_{1:t})\geq$ $\bigg{[}\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\bigg{[}\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\sum^{T-1}_{t=0}\Big{(}\log p_{\phi}(x_{t+1}|z_{t+1})$ (11)
$-D_{KL}(q_{\phi}(z_{t+1}|x_{t+1},z_{t},a_{t})\|p_{\phi}(z_{t+1}|z_{t},a_{t}))\Big{)}\bigg{]}\bigg{]}.$
For ease of notation, we have $q(z_{1}|x_{1},z_{0},a_{0}):=q(z_{1}|x_{1},s)$ and $p(z_{1}|z_{0},a_{0}):=p(z_{1})$. The ELBO provides a lower bound to the log-likelihood of the observed data, which is computationally intractable to compute directly. It is composed of two terms: the expected log-likelihood of the observed data given the latent variables, and the Kullback-Leibler (KL) divergence between the variational distribution and the prior distribution of the latent variables. Minimizing the KL divergence can be interpreted as enforcing the variational distribution to be as close as possible to the prior, while maximizing the expected log-likelihood ensures that the model accurately captures the distribution of the observed data. To predict the adherence, we have $\log p_{\phi}(y_{t+1}|z_{t+1})$ as a regularisor in the loss function.
$\min_{\phi}$ $\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\sum^{T-1}_{t=0}\left[D_{KL}\left(q_{\phi}(z_{t+1}|x_{t+1},z_{t},a_{t})\|p_{\phi}(z_{t+1}|z_{t},a_{t})\right)\right]\right]$ (12)
s.t. $\quad\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-1}_{t=0}-\log p_{\phi}(x_{t+1}|z_{t+1})\right]\right]\leq\xi_{\mathrm{score}}$ (13)
$\mathop{\mathbb{E}}_{(x_{1:T},a_{1:T-1},y_{1:T-1})\sim D}\left[\mathop{\mathbb{E}}_{z_{1:T}\sim q_{\phi}}\left[\sum^{T-2}_{t=0}-\log p_{\phi}(y_{t+1}|z_{t+1})\right]\right]\leq\xi_{\mathrm{adherence}}$ (14)
where in Eq. (13) we have regression with Gaussian distribution, and in Eq. (14) we use cross-binary entropy loss for classification. To solve the optimization problem, we incorporate the constraints into the objective function using Lagrange multipliers $\lambda$. We apply methods from (Chen et al., 2022) to adapt $\lambda$. This allows the model to balance the importance of the constraints relative to the divergence terms, which can help in avoiding common pitfalls in training such as suboptimal local minima and posterior collapse.

### v3

@@ Abstract
Objective: Subcutaneous Immunotherapy (SCIT) is the long-lasting causal treatment of allergic rhinitis (AR). How to enhance the adherence of patients to maximize the benefit of allergen immunotherapy (AIT) plays a crucial role in the management of AIT. This study aims to leverage novel machine learning models to precisely predict the risk of non-adherence of AR patients and related local symptom scores in three years SCIT.
Methods: The research develops and analyzes two models, sequential latent-variable model (SLVM) of Sequential Latent Actor-Critic (SLAC) and Long Short-Term Memory (LSTM) evaluating them based on scoring and adherence prediction capabilities.
Results: Excluding the biased samples at the first time step, the predictive adherence accuracy of the SLAC models is from 60% to 72%, and for LSTM models, it is 66% to 84%, varying according to the time steps. The range of Root Mean Square Error (RMSE) for SLAC models is between 0.93 and 2.22, while for LSTM models it is between 1.09 and 1.77. Notably, these RMSEs are significantly lower than the random prediction error of 4.55.
Conclusion: We creatively apply sequential models in the long-term management of SCIT with promising accuracy in the prediction of SCIT nonadherence in AR patients. While LSTM outperforms SLAC in adherence prediction, SLAC excels in score prediction for patients undergoing SCIT for AR. The state-action-based SLAC adds flexibility, presenting a novel and effective approach for managing long-term AIT.
@@ 2 Introduction
Allergic rhinitis (AR) is characterized by allergen-specific IgE-mediated inflammation in upper respiratory inflammation with a prevalence of up to 30% worldwide (Meltzer, 2016). In addition to allergen avoidance as the superior criterion, allergen-specific immunotherapy (AIT) aims to induce specific allergen immune tolerance, consequently achieving a status of clinical symptom remission. The repeatable intake of the specific unmodified or chemically modified allergens (allergoids) was the key to maintaining the symptoms. Among these approaches of AIT, subcutaneous immunotherapy (SCIT), sublingual immunotherapy (SLIT), and lymphatic immunotherapy (LIT) are demonstrated as the mainstream treatments regarding efficacy, safety, and side effects. Compared to the SLIT, SCIT is a clinic-dependent treatment in which the patient accepted an allergen extract injection subcutaneously. It is divided into the initial treatment stage (dose accumulation stage) and the maintenance treatment stage (dose maintenance stage). The World Allergy Organization (WAO) recommends that immunotherapy be maintained for three to five years and clinically recommended for at least two years. Patient adherence is a critical factor in ensuring long-lasting efficacy and sustaining symptom-relieving effects.
Due to the long duration of SCIT, cumbersome process, slow onset, individual differences in treatment effect, and other factors fundamentally impact the completeness of therapeutics. From the reported studies on AIT, the rate of adherence ranged from around 25% to over 90% (Passalacqua et al., 2013). The World health Organization (WHO) adopted the definition of “adherence” as “the extent to which a person’s behavior, such as taking medication, following a diet, or executing lifestyle changes, corresponds with agreed recommendations from a health care provider” (Organization et al., 2003). In recent European Academy of Allergy and Clinical Immunology (EAACI) guidelines, it is highlighted to educate patients on how immunotherapy works and on explaining the importance of compliance to the regular doses for three years of treatment (Roberts et al., 2018).
@@
In healthcare, machine learning, especially sequential models, stands at the forefront of innovation, providing new ways to analyze complex medical data and improve patient treatments. Previous research primarily concentrated on non-sequential prediction methods for adherence (Mousavi et al., 2022; Wang et al., 2020; Warren et al., 2022; Ruff et al., 2019). This approach presents a significant limitation in treatment processes, particularly for immunotherapy that often spans extended periods, such as three years. These non-sequential methods tend to predict only the overall outcome, overlooking the intricacies of intermediate time steps. To facilitate earlier intervention, a sequential model capable of making predictions at any given time step would be markedly more beneficial. While some subsequent studies have introduced sequential models (Hsu et al., 2022; Singh et al., 2022; Schleicher et al., 2023), their scope was restricted to predicting adherence alone. Our study enhances this approach by incorporating a state-action model, which can predict both adherence and score/state. This advancement allows for more precise and detailed management of AR patients by allergologists. These models excel in processing and analyzing time-dependent data, making them ideal for predicting patient adherence to treatments like SCIT for AIT. By effectively using sequential data, these algorithms uncover temporal patterns and correlations, leading to more accurate and personalized treatment plans.
In this study, in order to introduce the appropriate prediction model into long-period immunotherapy to customize the management of interventions and incorporate patient feedback, we have selected and evaluated two specific sequential models tailored to this scenario. Our findings demonstrate that these models are not only effective in predicting patient adherence to medical treatments but also invaluable in enhancing treatment strategies, thereby making a significant contribution to patient-centered healthcare.
@@ 3.1 Study design
The study design is a critical component that shapes the direction and reliability of our research. It includes a systematic approach to selecting the study population, the treatment methods applied, and the evaluation criteria (see Figure 1).
Figure 1: Flowchart on 205 patients treated with SCIT for Der p/Der f allergy during a 36-month treatment period. The adherence was assessed with sequential latent variable models focusing on patient’s demographic characteristics and clinical follow-up data.
@@ 3.1.1 Population
The study protocol was approved by the Ethics Committee of the First People’s Hospital of Foshan, Foshan, China. All methods were performed in accordance with the relevant guidelines and regulations.
@@ 3.1.2 SCIT treatment and evaluation
Before administering SCIT to enrolled patients, they will first perform a routine physical examination, inquire about related information since the last injection (including allergy symptoms), and post-injection, patients are observed for 30 minutes in case of the occurrence of side effects. Standardized adsorbed Der p and Der f allergen extracts (Allergopharma, Reinbeck, Germany) were used for SCIT. According to the manufacturer’s instructions, in the dose accumulation phase with weekly injections of allergen extracts with a gradually increased concentration from 100 SQ-U/mL to 10,000 SQ-U/mL, respectively injected 0.2, 0.4, 0.8 mL; after reaching the maintenance dose, 100,000 standardized quality units was used. In the maintenance phase, an injection interval of 6$\pm$2 weeks was carried out according to the manufacturer’s recommendations.
Patients receive regular treatment evaluations, including symptom scores and medication scores. The symptom score recorded a total of nasal symptoms (nasal itching, sneezing, rhinorrhea, nasal congestion), ocular symptoms (ocular itching, lacrimation), and pulmonary symptoms (shortness of breath, tightness in chest, perennial cough, wheezing), and assessed symptom severity using the visual analogue scale (VAS). In the VAS symptom score, the score of each symptom is from 0 to 10. 0 indicates that the patient has no discomfort and 10 indicates that the patient is extremely uncomfortable. The patient gives the score of each symptom according to the actual situation, and the sum of all symptom scores is the symptom score. Medication score recorded the use of current adjuvant medication within 1 month to reach symptom relief. The use of oral antihistamines, antileukotrienes, and bronchodilators were recorded as one point, local glucocorticoids as two points, oral glucocorticoids or combined medication (hormones and $\beta 2$ receptor agonists) as three points, and the total cumulative score was the medication score. Symptom scores and medication scores were assessed once at registration of SCIT and then thereafter.
@@ 3.1.3 Data Collection
Data were collected from patient records in hospitals, and the following information was extracted for analysis: patient age, gender, distance to clinic, ratio of AIT cost to family income, allergen test results, etc., as well as patient VAS system score and medication score information, including baseline data of patients before injection therapy, adverse reactions to SCIT. For the descriptive analysis, categorical variables were given as numbers and percentages, and continuous variables were presented using mean, standard deviation, median, interquartile range (IQR), and minimum and maximum values. To address missing values, we tracked every patient, which allowed us to ensure the dataset’s completeness. We did not remove outliers, aiming to follow real-clinical scenarios as closely as possible.
@@ 3.1.4 Survey methods
Adherence was defined as the accomplishment of three years of AIT. Non-adherence was defined as discontinuation of AIT at random time points during three years. The follow-up contents included (1) the main reasons for patients’ discontinuation of treatment; (2) the duration of discontinuation of treatment, and (3) Allergic symptoms after discontinuation of treatment.
@@ 3.2.1 Data
gender Female 62 (30.2) 22 40
Male 143 (69.8) 51 92
distance to clinic(km) $\leq 10$ 136 (66.3) 56 80
$>10$ 69 (33.7) 17 52
@@
Figure 2: Histogram of scores across six time steps. Score value (horizontal axis) vs. count (vertical axis).
@@
We have a dataset $D$, comprising sequences $x_{1},\dots,x_{T}\in\mathbb{R}^{11}$, $y_{1},\dots,y_{T-1}\in\mathbb{R}^{1}$, and a corresponding action $a_{t}\in R^{1}$. In the context of healthcare, the observations encompass $y_{t}$ (see Table 2) whether the patient will cease the treatment in the interval between the scoring measurements at $x_{t}$ and $x_{t+1}$. The actions $a_{t}$ represent the ongoing medical procedures for the patient during the period from $x_{t}$ to $x_{t+1}$ (see Fig. 2). In this context, $a_{t}$ is binary, reflecting whether treatment is given, and is numerically equivalent to the adherence variable $y_{t}$. Despite their numerical equivalence, we maintain a distinction between action $a_{t}$ and adherence $y_{t}$ to enhance model clarity and accommodate future research expansions, potentially allowing for a wider range of action values. For each patient, we possess basic information $s\in\mathbb{R}^{14}$ which includes age, gender, commute distance to clinic, ratio of cost to family income, eosinophils count, eosinophils percentage, nasal allergen provocation test (change of nasal resistance, $\Delta\text{NR}(\%)$), peak nasal inspiratory flow. $\Delta\mathrm{PNIF}(\%)$), serum total IgE level, sIgE of Dermatophagoides pteronyssinus (Derp), sIgE of Dermatophagoides farinae (Derf), skin prick test (Derp, Derf) (see Table 1 for more details).
3.2.2 Sequential latent variable model
@@
In our research, we use the Stochastic Latent Actor-Critic (SLAC) model (Lee et al., 2020). Our application differs from the original use of SLAC which is typically associated with reinforcement learning. Instead, we use its sequential latent-variable model (SLVM) without Actor-Critic. This approach aligns with similar methodologies found in other works (Krishnan et al., 2015; Karl et al., 2016; Gregor et al., 2018). The choice of the SLAC model was motivated by its ability to facilitate more efficient learning and superior generalization in intricate environments.
The SLVM is fundamentally a framework that processes information in a step-by-step manner, capturing the dynamics of an environment or process over time. It constructs a latent representation of the data that it identifies and uses underlying patterns or structures within the dataset that probably are not immediately obvious. This capability makes it exceptionally suitable for tasks where understanding temporal relationships is crucial, such as predicting patient adherence in allergen immunotherapy. The model operates by generating a sequence of predictions, each informed by the data received up to that point, thereby enabling it to adapt and refine its understanding as more information becomes available. This methodological choice allows our research to use the strengths of SLAC in a novel context, applying it to the predictive modeling of patient behaviors in a healthcare setting.
After training the model, given the historical data up to step $t-1$, the model is capable of generating a patient’s next scores from time step $t$ to $T$ directly from the latent space i.e., $p(x_{t:T}\mid x_{0:t-1},a_{0:t-1})$. Similarly, it can predict the adherence of step $t:T-1$ at time step $t$, $p(y_{t:T-1}\mid x_{0:t},a_{0:t-1})$. See details of the model and the training process in Appendix A.